# Menta: A Small Language Model for On-Device Mental Health Prediction

### v2

@@ 1 Introduction
Mental health disorders such as depression, anxiety, and suicidality affect hundreds of millions of people worldwide and constitute one of the leading contributors to the global burden of disease (world2017depression). Between 30% and 50% of people globally experience stress (piao2024continuous); an estimated 5.7% of adults suffer from depression (WHO2023depression); and more than 720,000 people die by suicide each year (WHO2025suicide). As mental health issues have continued to rise globally over the past few decades (goodwin2022trends; piao2024continuous; weaver2025global), there is an urgent need to enhance our understanding, diagnosis, and monitoring of these conditions through advanced technologies.
Despite growing demand for mental health support, significant diagnosis gaps persist globally. Barriers such as limited clinic hours, geographic inaccessibility, and workforce shortages hinder timely assessment and early self-awareness (WHOdiagnosis). Structural, patient-level, and systemic obstacles further delay diagnosis for those with serious mental illness (wiesepape2025behind). This mismatch between high need and limited resources reflects the constraints of traditional diagnostic models (stein2022psychiatric). Moreover, these traditional delivery of in-person sessions and self-reported questionnaires struggles to scale, and faces systemic bottlenecks, including workforce shortages (WHO_SIMH; brahmbhatt2024access), high costs (patel2018lancet), geographic disparities (yang2025wait), and stigma (clement2015impact). Mental health wait times can range from 3 to 18 months, varying by region and provider availability (mcmahan2022using; yang2025wait; patel2018lancet). With the growing need for early detection and real-time monitoring, developing more seamless methods to identify early signs is critical for delivering timely support (patel2018lancet). This underscores the importance of portable, resource-efficient solutions such as small language models for scalable mental health detection and intervention.
@@
The rise of portable, on‑device technologies offers a promising path for the early detection of mental health issues, particularly through passive, real‑time analysis of social media posts and smartphone applications. Smartphone-based linguistic data from social media enables scalable, passive monitoring of psychological states, facilitating early identification of mood disturbances, stress, or suicide risk. For example, such systems can support suicide prevention efforts (arean2016mobile) and capture patterns in social media use that are relevant to mental well-being (hamilton2025improving). Running locally on personal devices, these models preserve user privacy while analyzing language inputs for indicators of depression or stress (shin2023fedtherapist). As a result, on-device deployment enables scalable, privacy-preserving mental health monitoring, extending detection beyond clinical settings and making access easier for a broader population.
In this context, language models offer a powerful means of addressing the limitations of traditional approaches to mental health monitoring (xu2024mental). Language models outperform supervised learning classifiers in mental health prediction tasks for capturing deeper context, information acquisition, and generalization capabilities (jin2025applications), making them more practical for web-scale and real-world mental health monitoring.
However, the application of large language models (LLMs) faces significant challenges: they are computationally expensive to train and deploy, and they raise privacy concerns in terms of processing sensitive data on remote servers (jin2025applications). Their large size also makes them impractical for resource-constrained on-device settings, underscoring the need for more efficient alternatives. By contrast, Small Language Models (SLMs), offer comparable or even superior potential for digital health applications due to their lightweight architectures, lower inference costs, and feasibility of local deployment (abdin2024phi; jia2025beyond). Deploying on digital devices locally, a privacy-preserving, on-device workflow is expected for mental health prediction from social media activity using lightweight SLMs which emphasizes in-situ AI screening across multiple mental health dimensions without storing personal data, enabling real-time, low-resource, and ethically-aligned monitoring (Figure 1). Yet, the potential of SLMs for mental health prediction remains largely underexplored.
Despite work on the development of SLMs, there is no agreement yet on a standardized definition of SLMs (wang2024comprehensive). Given the lightweight requirement for such real-time mental health status detection tasks on social media posts, we define SLMs to be models ranging from a minimum of 1B to strictly under 7B parameters are commonly considered ‘small’ in contrast to full-scale LLMs (10B+) while those more than 7B to be large-scale small language models (LSSLMs) (abdin2024phi; team2024qwen2).
@@ 2 Related Work
Social Media Posts: Indication of Mental Health. Social media posts offer a valuable lens into individuals’ emotions, thoughts, and mental health status. Globally, approximately 4.8 billion people, nearly 60% of the population, use social media, spending a combined 11.5 billion hours on these platforms each day UMaine2023. With the increasing use of social media platforms for emotional expression and peer support, user-generated text has become an important resource for understanding psychological well-being (hussain2025optimised; zhunis2022emotion) and identifying early warning signs of mental distress (chancellor2020methods; de2013predicting). For instance, a study (kim2020deep) built a classifier on Reddit posts to detect multiple mental disorders, while the EmoMent corpus (atapattu2022emoment) shows that emotion annotations in user text correlate with mental illness indicators even in non-Western populations. These findings support the potential for social media–based models to serve as early warning systems for mental health support.
@@
Early approaches for textual information analysis among social media posts relied on traditional machine learning and feature-engineered models to classify or detect mental health status. For example, Jiang et al. used contextualized word embeddings such as ELMo and BERT to detect mental health conditions from Reddit posts (jiang2020detection), showing that deep contextualized models significantly outperform traditional bag-of-words or static embeddings for mental health classification tasks. Similarly, Sarkar et al. developed a multi-task learning model that jointly predicts depression and anxiety from Reddit posts, which outperforms single-task baselines by leveraging shared features across related conditions (sarkar2022predicting). Another studies developed a scalable deep-learning screening tool for suicide risk with potential clinical implications for early intervention such as CNN (coppersmith2018natural), LSTM (coppersmith2018natural) and SVM (ji2018supervised).
@@
More recently, LLMs have been adapted to this domain. A prior study used ChatGPT in zero‑shot classification tasks across three mental health domains (stress, depression, suicidality) on social media datasets, validating the ability of LLMs with no prior knowledge in mental health prediction tasks (lamichhane2023evaluation). Another study focuses only on the depression detection task from online text while collaborates the finding that fine-tuned LSSLMs achieve great improvements over prior state‑of‑art models (shah2025advancing). Moreover, the interpretability of LLMs in prediction tasks such as depressive symptoms (bolegave2025gold; chen2025generating; belcastro2025detecting) and emotional states (yang2023towards) have also been explored, with the finding that prompt engineering with emotional cues and few‑shot examples improves performance. Importantly, compared to traditional algorithmic models, lagnuage models offer not only strong predictive capabilities but also human-readable interpretations and the flexibility to support multiple downstream tasks within a unified framework.
@@
A more comprehensive study introduced Mental-LLM (xu2024mental), demonstrated that instruction-tuned LLMs show large gains in balanced accuracy and can outperform task-specific baselines, with the best finetuned models (Mental‑Alpaca and Mental‑FLAN‑T5) outperform GPT‑3.5 by approximately 10.9% and beat GPT‑4 by about 4.8%. They also improved model generalizability fine-tuned on diverse datasets. Follow-up work has extended this line of research with domain-specialized models. Similarly, Shi et al. proposed a LSSLM called MentalQLM for mental health tasks by using a base backbone (7B) and then applying a dual LoRA strategy (hu2022lora), achieving comparable or even superior performance to much larger LLMs on several mental health diagnostic tasks (shi2025mentalqlm). While these studies demonstrate the feasibility of using LLMs for mental health outcome prediction, significant challenges persist, including concerns over privacy and security (sarwar2025fedmentalcare), deployment limitations (maurya2025exploring), and high computational costs (wang2023chatgpt).
@@
Small Language Models and Mental Health. To address this gap of application bottlenecks, SLMs emerged as a pragmatic alternative of LLMs (van2024survey), optimized for efficiency through techniques such as quantization (frantar2022gptq), knowledge distillation (kang2023knowledge) and light‑weight architectures (wang2025survey). SLMs were designed to deliver core Natural Language Processing capabilities such as text classification (pecher2024comparing), summarization (wang2025distilling), and question‑answering (lee2024can), in resource‑constrained settings including mobile devices, embedded systems, and edge computing.
@@
Offering a compelling alternative to LLMs, SLMs reduce computational overhead while retaining strong performance in many specialized tasks. Moreover, recent work shows that in classification settings with limited data, well‑tuned smaller models can match or even surpass larger models when the task is focused and domain‑specific including text classification (lepagnol2024small; luo2023exploring) and healthcare domains (gondara2025small). These findings highlight that SLMs offer an efficient, scalable, and domain‑optimized solution, making them a promising direction for mental‑health detection from social media. However, previous work has typically focused on a single task or domain and has not explored multi-domain settings that combine multiple mental health conditions such as depression, stress, and suicidality.
@@
Unlike prior LLM-based (sarkar2022predicting; xu2024mental; yang2024mentallama; kim2024mhgpt; shi2025mentalqlm) which range from 7B to 70B parameters which are impractical for deplyment in resource-constrained or privacy-sensitive settings, lightweight SLMs emphasize deployability, reporting inference latency, memory footprint, and approximate cost efficiency. Prior work evaluated several SLMs against LLM baselines across multiple individual mental health understanding tasks with zero-shot and few-shot prompting trained on social media datasets, and concluded that few-shot prompting helps SLMs more (jia2025beyond). Also, Kim et al. developed a SLM called mhGPT, which outperforms larger models such as MentaLLaMA and Gemma in mental health tasks despite having far fewer parameters and using less data (kim2024mhgpt). A prior study also suggests SLMs suitable to be integrated in clinical workflows for structured and clinically meaningful tasks beyond generic detection (aich2024using). However, these studies lack systematic task balancing, rarely optimized with parameter-efficient fine-tuning (LoRA), meaning adaptation is either minimal (prompting) or heavy (full fine-tuning).
@@
Deployment and On-Device Language Models. On‑device AI offers a compelling opportunity for mental health monitoring by enabling personalized assessment in real time, while preserving user privacy, reducing latency, and expanding accessibility. Running inference directly on a user’s smartphone avoids sending sensitive linguistic and behavioural data to cloud servers, thereby lowering risks of data leakage and minimizing network dependencies (mandal2025towards). Furthermore, mobile and edge deployment largely improves accessibility by bringing mental health support tools directly to users, addressing traditional barriers such as geographic inaccessibility and long wait times for in‑clinic screening (bunyi2021accessibility). Additionally, because on‑device models eliminate round‑trip communication with remote servers, they deliver near‑instant responses and support real‑time detection of changes in language or behaviour indicative of stress or mood deterioration (ni2025scoping).
@@
Mobile and edge-based mental health monitoring using smartphones and wearable sensors has gained momentum, offering significant advantages in latency, energy efficiency, and data privacy over cloud-dependent systems. Recent work has demonstrated effective deployment of SLMs on consumer devices, such as the Samsung Galaxy S24, for tasks like document assistance (pham2024slimlm) and health prediction (wang2025healthslm), with substantial reductions in memory usage and inference latency. Conceptual analyses further argue that SLMs are better suited than LLMs for real-world, interactive applications due to their lightweight nature and ease of deployment under resource constraints (belcak2025small). On-device or near-device processing eliminates the need for continuous cloud communication, thereby enabling real-time detection while safeguarding user data—an essential consideration for privacy (lu2024small).
@@ 4.1 Datasets and Task Definitions
Dreaddit turcan2019dreaddit: From a multi-domain stress corpus, 3,500 segments were manually annotated via Amazon Mechanical Turk as either “stress” or “not stress”, ensuring coverage across diverse domains.
DepSeverity naseem2022early: This dataset is aimed at accurately identifying users’ depression severity levels, annotated by CLEF eRisk organizers based on DSAS and clinical standards, categorizing the dataset into four severity levels: Minimal, Mild, Moderate, and Severe.
SDCNL haque2021deep: This dataset of 1,895 Reddit posts was collected from communities related to depression and suicidal ideation. Posts from suicide-related communities were labeled as “suicidal ideation,” and posts from depression communities were labeled as ”non-suicidal ideation.”, subsequently employed unsupervised labeling methods to correct potential misclassifications.
CSSRS-Suicide gaur2019knowledge: This dataset contains samples related to depressive suicidal ideation and behaviors. Five hundred users were randomly selected, and domain experts annotated the data using the Columbia-Suicide Severity Rating Scale (brown2020c). The resulting labels fall into five categories: supportive, indicator, ideation, behavior, and attempt.
@@ 4.2 Baseline Experiments
In this study, we evaluate a diverse set of SLMs limited to under 7B parameters. Phi-3 (3.8B) and Phi-3.5 (3.8B) are lightweight Microsoft models designed for reasoning and efficiency (abdin2024phi). LLaMA 3.2 (3B) is Meta’s compact release optimized for low-resource deployment (dubey2024llama). Gemma-3 (1B and 4B) are Google’s instruction-tuned models emphasizing accessibility and safety (team2025gemma). Qwen-2.5 (3B) and Qwen-3 (4B) are Alibaba’s models tailored for multilingual and domain adaptability (yang2025qwen3). Phi-4 Mini (3.8B) extends Microsoft’s efficiency line (abouelenin2025phi). TinyLLaMA (1.1B) (zhang2024tinyllama) and Falcon (1.3B) (almazrouei2023falcon) are streamlined open-source transformer models for edge deployment. StableLM (3B) is Stability AI’s open family prioritizing lightweight deployment and transparency (pinnaparaju2024stable).
We developed a multi-task fine-tuned model called Menta employed a shared Qwen3-4B backbone with LoRA adapters, jointly trained on all six tasks. Training was balanced using task-specific sampling weights, and the proposed combined loss function (§3.2) was applied to each task. We compare the developed Menta model using the adjusted loss function with other selected fine-tuned SLMs, including Phi-4 Mini (3.8B), StableLM (3B) and Falcon-1.3B. In addition, we compare the SLMs against larger-scale baselines, Mental-Alpaca and Mental-FLAN-T5 xu2024mental, developed by Xu et al. through multi-task instruction fine-tuning of Alpaca (7B) taori2023stanford and FLAN-T5 (11B) chung2024scaling using the same datasets and tasks.

### v3

@@ 1 Introduction
Mental health disorders such as depression, anxiety, and suicidality affect hundreds of millions of people worldwide and constitute one of the leading contributors to the global burden of disease (Organization et al., 2017). Between 30% and 50% of people globally experience stress (Piao et al., 2024); an estimated 5.7% of adults suffer from depression (World Health Organization, 2023); and more than 720,000 people die by suicide each year (World Health Organization, 2025a). As mental health issues have continued to rise globally over the past few decades (Goodwin et al., 2022; Piao et al., 2024; Weaver et al., 2025), there is an urgent need to enhance our understanding, diagnosis, and monitoring of these conditions through advanced technologies.
Despite growing demand for mental health support, significant diagnosis gaps persist globally. Barriers such as limited clinic hours, geographic inaccessibility, and workforce shortages hinder timely assessment and early self-awareness (World Health Organization, 2025b). Structural, patient-level, and systemic obstacles further delay diagnosis for those with serious mental illness (Wiesepape et al., 2025). This mismatch between high need and limited resources reflects the constraints of traditional diagnostic models (Stein et al., 2022). Moreover, these traditional delivery of in-person sessions and self-reported questionnaires struggles to scale, and faces systemic bottlenecks, including workforce shortages (Organization, 2025; Brahmbhatt & Schpero, 2024), high costs (Patel et al., 2018), geographic disparities (Yang & Zhang, 2025), and stigma (Clement et al., 2015). Mental health wait times can range from 3 to 18 months, varying by region and provider availability (McMahan et al., 2022; Yang & Zhang, 2025; Patel et al., 2018). With the growing need for early detection and real-time monitoring, developing more seamless methods to identify early signs is critical for delivering timely support (Patel et al., 2018). This underscores the importance of portable, resource-efficient solutions such as small language models for scalable mental health detection and intervention.
@@
The rise of portable, on‑device technologies offers a promising path for the early detection of mental health issues, particularly through passive, real‑time analysis of social media posts and smartphone applications. Smartphone-based linguistic data from social media enables scalable, passive monitoring of psychological states, facilitating early identification of mood disturbances, stress, or suicide risk. For example, such systems can support suicide prevention efforts (Areán et al., 2016) and capture patterns in social media use that are relevant to mental well-being (Hamilton et al., 2025). Running locally on personal devices, these models preserve user privacy while analyzing language inputs for indicators of depression or stress (Shin et al., 2023). As a result, on-device deployment enables scalable, privacy-preserving mental health monitoring, extending detection beyond clinical settings and making access easier for a broader population.
In this context, language models offer a powerful means of addressing the limitations of traditional approaches to mental health monitoring (Xu et al., 2024). Language models outperform supervised learning classifiers in mental health prediction tasks for capturing deeper context, information acquisition, and generalization capabilities (Jin et al., 2025), making them more practical for web-scale and real-world mental health monitoring.
However, the application of large language models (LLMs) faces significant challenges: they are computationally expensive to train and deploy, and they raise privacy concerns in terms of processing sensitive data on remote servers (Jin et al., 2025). Their large size also makes them impractical for resource-constrained on-device settings, underscoring the need for more efficient alternatives. By contrast, Small Language Models (SLMs), offer comparable or even superior potential for digital health applications due to their lightweight architectures, lower inference costs, and feasibility of local deployment (Abdin et al., 2024; Jia et al., 2025). Deploying on digital devices locally, a privacy-preserving, on-device workflow is expected for mental health prediction from social media activity using lightweight SLMs which emphasizes in-situ AI screening across multiple mental health dimensions without storing personal data, enabling real-time, low-resource, and ethically-aligned monitoring (Figure 1). Yet, the potential of SLMs for mental health prediction remains largely underexplored.
Despite work on the development of SLMs, there is no agreement yet on a standardized definition of SLMs (Wang et al., 2024). Given the lightweight requirement for such real-time mental health status detection tasks on social media posts, we define SLMs to be models ranging from a minimum of 1B to strictly under 7B parameters are commonly considered ‘small’ in contrast to full-scale LLMs (10B+) while those more than 7B to be large-scale small language models (LSSLMs) (Abdin et al., 2024; Team, 2024).
@@ 2 Related Work
Social Media Posts: Indication of Mental Health. Social media posts offer a valuable lens into individuals’ emotions, thoughts, and mental health status. Globally, approximately 4.8 billion people, nearly 60% of the population, use social media, spending a combined 11.5 billion hours on these platforms each day University of Maine (2023). With the increasing use of social media platforms for emotional expression and peer support, user-generated text has become an important resource for understanding psychological well-being (Hussain et al., 2025; Zhunis et al., 2022) and identifying early warning signs of mental distress (Chancellor & De Choudhury, 2020; De Choudhury et al., 2013). For instance, a study (Kim et al., 2020) built a classifier on Reddit posts to detect multiple mental disorders, while the EmoMent corpus (Atapattu et al., 2022) shows that emotion annotations in user text correlate with mental illness indicators even in non-Western populations. These findings support the potential for social media–based models to serve as early warning systems for mental health support.
@@
Early approaches for textual information analysis among social media posts relied on traditional machine learning and feature-engineered models to classify or detect mental health status. For example, Jiang et al. used contextualized word embeddings such as ELMo and BERT to detect mental health conditions from Reddit posts (Jiang et al., 2020), showing that deep contextualized models significantly outperform traditional bag-of-words or static embeddings for mental health classification tasks. Similarly, Sarkar et al. developed a multi-task learning model that jointly predicts depression and anxiety from Reddit posts, which outperforms single-task baselines by leveraging shared features across related conditions (Sarkar et al., 2022). Another studies developed a scalable deep-learning screening tool for suicide risk with potential clinical implications for early intervention such as CNN (Coppersmith et al., 2018), LSTM (Coppersmith et al., 2018) and SVM (Ji et al., 2018).
@@
More recently, LLMs have been adapted to this domain. A prior study used ChatGPT in zero‑shot classification tasks across three mental health domains (stress, depression, suicidality) on social media datasets, validating the ability of LLMs with no prior knowledge in mental health prediction tasks (Lamichhane, 2023). Another study focuses only on the depression detection task from online text while collaborates the finding that fine-tuned LSSLMs achieve great improvements over prior state‑of‑art models (Shah et al., 2025). Moreover, the interpretability of LLMs in prediction tasks such as depressive symptoms (Bolegave & Bhattacharya, 2025; Chen & Lin, 2025; Belcastro et al., 2025) and emotional states (Yang et al., 2023) have also been explored, with the finding that prompt engineering with emotional cues and few‑shot examples improves performance. Importantly, compared to traditional algorithmic models, lagnuage models offer not only strong predictive capabilities but also human-readable interpretations and the flexibility to support multiple downstream tasks within a unified framework.
@@
A more comprehensive study introduced Mental-LLM (Xu et al., 2024), demonstrated that instruction-tuned LLMs show large gains in balanced accuracy and can outperform task-specific baselines, with the best finetuned models (Mental‑Alpaca and Mental‑FLAN‑T5) outperform GPT‑3.5 by approximately 10.9% and beat GPT‑4 by about 4.8%. They also improved model generalizability fine-tuned on diverse datasets. Follow-up work has extended this line of research with domain-specialized models. Similarly, Shi et al. proposed a LSSLM called MentalQLM for mental health tasks by using a base backbone (7B) and then applying a dual LoRA strategy (Hu et al., 2022), achieving comparable or even superior performance to much larger LLMs on several mental health diagnostic tasks (Shi et al., 2025). While these studies demonstrate the feasibility of using LLMs for mental health outcome prediction, significant challenges persist, including concerns over privacy and security (Sarwar, 2025), deployment limitations (Maurya et al., 2025), and high computational costs (Wang et al., 2023).
@@
Small Language Models and Mental Health. To address this gap of application bottlenecks, SLMs emerged as a pragmatic alternative of LLMs (Van Nguyen et al., 2024), optimized for efficiency through techniques such as quantization (Frantar et al., 2022), knowledge distillation (Kang et al., 2023) and light‑weight architectures (Wang et al., 2025a). SLMs were designed to deliver core Natural Language Processing capabilities such as text classification (Pecher et al., 2024), summarization (Wang et al., 2025b), and question‑answering (Lee et al., 2024), in resource‑constrained settings including mobile devices, embedded systems, and edge computing.
@@
Offering a compelling alternative to LLMs, SLMs reduce computational overhead while retaining strong performance in many specialized tasks. Moreover, recent work shows that in classification settings with limited data, well‑tuned smaller models can match or even surpass larger models when the task is focused and domain‑specific including text classification (Lepagnol et al., 2024; Luo et al., 2023) and healthcare domains (Gondara et al., 2025). These findings highlight that SLMs offer an efficient, scalable, and domain‑optimized solution, making them a promising direction for mental‑health detection from social media. However, previous work has typically focused on a single task or domain and has not explored multi-domain settings that combine multiple mental health conditions such as depression, stress, and suicidality.
@@
Unlike prior LLM-based (Sarkar et al., 2022; Xu et al., 2024; Yang et al., 2024b; Kim et al., 2024; Shi et al., 2025) which range from 7B to 70B parameters which are impractical for deplyment in resource-constrained or privacy-sensitive settings, lightweight SLMs emphasize deployability, reporting inference latency, memory footprint, and approximate cost efficiency. Prior work evaluated several SLMs against LLM baselines across multiple individual mental health understanding tasks with zero-shot and few-shot prompting trained on social media datasets, and concluded that few-shot prompting helps SLMs more (Jia et al., 2025). Also, Kim et al. developed a SLM called mhGPT, which outperforms larger models such as MentaLLaMA and Gemma in mental health tasks despite having far fewer parameters and using less data (Kim et al., 2024). A prior study also suggests SLMs suitable to be integrated in clinical workflows for structured and clinically meaningful tasks beyond generic detection (Aich et al., 2024). However, these studies lack systematic task balancing, rarely optimized with parameter-efficient fine-tuning (LoRA), meaning adaptation is either minimal (prompting) or heavy (full fine-tuning).
@@
Deployment and On-Device Language Models. On‑device AI offers a compelling opportunity for mental health monitoring by enabling personalized assessment in real time, while preserving user privacy, reducing latency, and expanding accessibility. Running inference directly on a user’s smartphone avoids sending sensitive linguistic and behavioural data to cloud servers, thereby lowering risks of data leakage and minimizing network dependencies (Mandal et al., 2025). Furthermore, mobile and edge deployment largely improves accessibility by bringing mental health support tools directly to users, addressing traditional barriers such as geographic inaccessibility and long wait times for in‑clinic screening (Bunyi et al., 2021). Additionally, because on‑device models eliminate round‑trip communication with remote servers, they deliver near‑instant responses and support real‑time detection of changes in language or behaviour indicative of stress or mood deterioration (Ni & Jia, 2025).
@@
Mobile and edge-based mental health monitoring using smartphones and wearable sensors has gained momentum, offering significant advantages in latency, energy efficiency, and data privacy over cloud-dependent systems. Recent work has demonstrated effective deployment of SLMs on consumer devices, such as the Samsung Galaxy S24, for tasks like document assistance (Pham et al., 2024) and health prediction (Wang et al., 2025c), with substantial reductions in memory usage and inference latency. Conceptual analyses further argue that SLMs are better suited than LLMs for real-world, interactive applications due to their lightweight nature and ease of deployment under resource constraints (Belcak et al., 2025). On-device or near-device processing eliminates the need for continuous cloud communication, thereby enabling real-time detection while safeguarding user data—an essential consideration for privacy (Lu et al., 2024).
@@ 4.1 Datasets and Task Definitions
Dreaddit Turcan & McKeown (2019): From a multi-domain stress corpus, 3,500 segments were manually annotated via Amazon Mechanical Turk as either “stress” or “not stress”, ensuring coverage across diverse domains.
DepSeverity Naseem et al. (2022): This dataset is aimed at accurately identifying users’ depression severity levels, annotated by CLEF eRisk organizers based on DSAS and clinical standards, categorizing the dataset into four severity levels: Minimal, Mild, Moderate, and Severe.
SDCNL Haque et al. (2021): This dataset of 1,895 Reddit posts was collected from communities related to depression and suicidal ideation. Posts from suicide-related communities were labeled as “suicidal ideation,” and posts from depression communities were labeled as ”non-suicidal ideation.”, subsequently employed unsupervised labeling methods to correct potential misclassifications.
CSSRS-Suicide Gaur et al. (2019): This dataset contains samples related to depressive suicidal ideation and behaviors. Five hundred users were randomly selected, and domain experts annotated the data using the Columbia-Suicide Severity Rating Scale (Brown et al., 2020). The resulting labels fall into five categories: supportive, indicator, ideation, behavior, and attempt.
@@ 4.2 Baseline Experiments
In this study, we evaluate a diverse set of SLMs limited to under 7B parameters. Phi-3 (3.8B) and Phi-3.5 (3.8B) are lightweight Microsoft models designed for reasoning and efficiency (Abdin et al., 2024). LLaMA 3.2 (3B) is Meta’s compact release optimized for low-resource deployment (Dubey et al., 2024). Gemma-3 (1B and 4B) are Google’s instruction-tuned models emphasizing accessibility and safety (Team et al., 2025). Qwen-2.5 (3B) and Qwen-3 (4B) are Alibaba’s models tailored for multilingual and domain adaptability (Yang et al., 2025). Phi-4 Mini (3.8B) extends Microsoft’s efficiency line (Abouelenin et al., 2025). TinyLLaMA (1.1B) (Zhang et al., 2024) and Falcon (1.3B) (Almazrouei et al., 2023) are streamlined open-source transformer models for edge deployment. StableLM (3B) is Stability AI’s open family prioritizing lightweight deployment and transparency (Pinnaparaju et al., 2024).
We developed a multi-task fine-tuned model called Menta employed a shared Qwen3-4B backbone with LoRA adapters, jointly trained on all six tasks. Training was balanced using task-specific sampling weights, and the proposed combined loss function (§3.2) was applied to each task. We compare the developed Menta model using the adjusted loss function with other selected fine-tuned SLMs, including Phi-4 Mini (3.8B), StableLM (3B) and Falcon-1.3B. In addition, we compare the SLMs against larger-scale baselines, Mental-Alpaca and Mental-FLAN-T5 Xu et al. (2024), developed by Xu et al. through multi-task instruction fine-tuning of Alpaca (7B) Taori et al. (2023) and FLAN-T5 (11B) Chung et al. (2024) using the same datasets and tasks.